OpenStudy (anonymous):

solve each literal equation for the given variable Solve S = 4v2 for v.(1 point) Solve M = 2x + 3y for y. (1 point) Solve t = the quantity p plus 3 r, divided by six for r. (1 point) Solve V = π r2h for h. (1 point) Solve P = 2(l + w) for l. What are the missing values in the table? (1 point)

OpenStudy (anonymous):

SHOW ALL WORK

OpenStudy (anonymous):

\[S=4v^2\] \[\frac{S}{4}=v^2\] \[\pm\sqrt{\frac{S}{4}}=v\] \[v=\pm\frac{\sqrt{S}}{2}\] is the first one

OpenStudy (anonymous):

\[M = 2x + 3y\] \[M-2x=3y\] \[y=\frac{M-2x}{3}\] for the next you good from there?

OpenStudy (anonymous):

NOOO LOL

OpenStudy (anonymous):

\[t=\frac{p+3r}{6}\] \[6t=p+3r\] \[6t-p=3r\] \[r=\frac{6t-p}{3}\]

OpenStudy (anonymous):

\[V=\pi r^2 h\] \[h=\frac{V}{\pi r^2}\]

OpenStudy (anonymous):

P w l 14 2 22 8 ?

OpenStudy (anonymous):

thats the table

OpenStudy (anonymous):

and last one is \[\frac{P}{2}=l+w\] \[\frac{P}{2}-w=l\]

OpenStudy (anonymous):

by the answer above \[l=\frac{P}{2}-w\] so if \[P=22, w = 8\] you get \[l=\frac{22}{2}-8=11-8=3\]